# On continuity of state-dependent utilities

### v1

@@ Appendix A Appendix: auxiliary results
Lemma A.1
@@
Proof 5
We prove Item (i) in the spirit of KS91 (Proposition 1.1.13 and Remark 1.1.14) as follows. Let $m>0$. For $n\geq 1$, $k=0,1,...,2\cdot 2^{n}-1$, we define
@@
[…]
Proposition A.2
@@
[…]
Proof 6
In the present proof we shall rely on few concepts related to the theory of stochastic processes and in particular on stopping times. We refer the reader to KS91 , page 6, for an exhaustive treatment of the notions involved. Lemma A.1 guarantees that $T_{u}:\mathcal{L}^{\infty}(\Omega,{\mathcal{F}})\to\mathbb{R}$ is well defined and finite valued. Fix $M>0$ and define $X_{t}(\omega):=u(\omega,t-M)$ for $t\geq 0$, omitting the dependence on $M$ for simplicity. Take also the filtration ${\mathcal{F}}_{t}:={\mathcal{F}},t\geq 0$, which is trivially right continuous by definition: $\bigcap_{s>t}{\mathcal{F}}_{s}={\mathcal{F}}={\mathcal{F}}_{t}$. Then $(X_{t})_{t}$ is an RCLL process, adapted to $({\mathcal{F}}_{t})_{t}$. Set $\Delta X_{0}=0$ and for $t>0$ set $\Delta X_{t}:=X_{t}-\sup_{0\leq s<t}X_{s}=X_{t}-\lim_{s\uparrow t}X_{s}$. Following the idea in So13 , we define $\tau^{\varepsilon}$ as the first jump time for the size $\varepsilon>0$, namely
@@
where equality ()normal-⋆($\star$)( ⋆ ) follows from posing $\Delta X_{0}=0$. Looking back at $u(\omega,\cdot)$, $\tau^{\varepsilon}(\omega)$ is the first time $u(\omega,\cdot)$ jumps, in the interval $(-M,+\infty)$, with a jump size strictly greater than $\varepsilon$, with the usual convention that such time is set to be +$\infty$+ ∞ if no such jump actually occurs. By So13 , $\tau^{\varepsilon}:\Omega\rightarrow[0,+\infty]$ is a $({\mathcal{F}}_{t})_{t}$-stopping time, hence in particular it is ${\mathcal{F}}$-measurable by definition of such a filtration. Setting
@@
Take now $s_{k}:=(Z-\frac{1}{k}\mathbf{1}_{B})^{+}$ for $k\in\mathbb{N}$. $s_{k}$ is ${\mathcal{F}}-\mathbb{B}_{[0,+\infty)}$ measurable and bounded. Observe that since $(X_{t})_{t}$ is RCLL, by KS91 Remark 1.1.14 we have $(t,\omega)\mapsto X_{t}(\omega)$ is $\mathbb{B}_{[0,+\infty)}\otimes{\mathcal{F}}-\mathbb{B}_{[0,+\infty)}$ measurable, and $\omega\mapsto(\omega,Y(\omega))$ is ${\mathcal{F}}-{\mathcal{F}}\otimes\mathbb{B}_{[0,+\infty)}$ measurable for every ${\mathcal{F}}-\mathbb{B}_{[0,+\infty)}$ measurable bounded $Y:\Omega\rightarrow[0,+\infty)$. Hence $\omega\mapsto X_{s_{k}(\omega)}(\omega)$ is ${\mathcal{F}}-\mathbb{B}_{[0,+\infty)}$ measurable. By continuity from below, argued at the beginning of this proof,
@@
Corollary A.3
Under the same assumptions of Proposition A.2, there exists a function $\widehat{u}$ such that $\widehat{u}(\omega,\cdot)$ is nondecreasing and continuous for every $\omega\in\Omega$, $E=\{\omega\in\Omega\mid u(\omega,x)=\widehat{u}(\omega,x)\,\forall x\in\mathbb{R}\}\in{\mathcal{F}}$ and ${\mathbb{P}}(E)=1$. In particular, $T_{u}(f)=T_{\widehat{u}}(f),\,\forall f\in\mathcal{L}^{\infty}(\Omega,{\mathcal{F}})$. Finally, if $u(\omega,\cdot)$ is strictly increasing on $\mathbb{R}$ for every $\omega\in\Omega$, $\widehat{u}(\omega,\cdot)$ can be taken strictly increasing on $\mathbb{R}$ for every $\omega\in\Omega$.
Proof 7
As a consequence of Proposition A.2, $A_{cont}\in{\mathcal{F}}$. Set $A:=A_{cont}$ and define now $\widehat{u}(\omega,x):=u(\omega,x)\mathbf{1}_{A}(\omega)+x\mathbf{1}_{\Omega\setminus A}$. Then clearly $\widehat{u}(\omega,\cdot)$ is continuous and nondecreasing either $\omega\in A$ or $\omega\in\Omega\setminus A$. Further for any fixed $x\in\mathbb{R}$ the map $\omega\mapsto\widehat{u}(\omega,x)$ is ${\mathcal{F}}$ measurable. As $u(\omega,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ implies $\widehat{u}(\omega,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ for every $x\in\mathbb{R}$ we can invoke Lemma A.1 to obtain that $\widehat{u}$ is ${\mathcal{F}}\otimes\mathbb{B}_{\mathbb{R}}-\mathbb{B}_{\mathbb{R}}$ measurable, $T_{\widehat{u}}$ is well defined in ${\mathcal{L}}^{\infty}(\Omega,{\mathcal{F}})$ and $T_{u}\equiv T_{\widehat{u}}$.
Ethics declarations
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v2

@@ 5.3 Proof of Theorem 8
5.4 Proof of Corollary 5
The proof of the corollary is merely an application of the results in the following sections, we briefly outline here the main steps. To show that (i) implies (iv), consider a preference order $\succeq$ that satisfies (SM), (PC) and (ST) and suppose ${\mathcal{F}}$ contains at least three disjoint non-null events. The existence of a conditional Chisini mean under these assumptions is guaranteed by [11], Theorem 3.2 (See the proof of Proposition 9 for more details). Representability of conditional Chisini means is then obtained from Theorem 7 and Theorem 8 as follows. The former provides a pair ($\mathbb{u},\mathbb{P}$) such that $\mathbb{u}:\Omega\times\mathbb{R}\to\mathbb{R}$ is ${\mathcal{F}}$-regular and $\mathbb{P}$-integrable. The latter allows the construction for any ${\mathcal{G}}\subseteq{\mathcal{F}}$ of a ${\mathcal{G}}$-regular function $\mathbb{u}_{\mathcal{G}}:\Omega\times\mathbb{R}\to\mathbb{R}$ that is $\mathbb{P}$-integrable and such that, for all $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ and for a suitable version $h$ of $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,f)|{\mathcal{G}}\right]$, we have $\mathbb{u}_{\mathcal{G}}^{-1}\left(\cdot,h(\cdot)\right)\in\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$.
Regarding the converse implication, observe that Item (iv) actually implies that any representing functional $T$ for $\succeq$ is conditionable for every choice of ${\mathcal{G}}\subseteq{\mathcal{F}}$. Therefore, the same argument used in the proof of Theorem 4 (in particular Item (iii) $\Rightarrow$ Item (i)) shows that $\succeq$ satisfies (ST).
5.5 Proof of Theorem 2
In this section we deal with equivalence classes of random variables, and therefore for any $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ we denote by $X=[f]_{\mathbb{P}}$ the element in $L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$ such that $f\in X$.
We first prove the converse implication of the statement. To this aim consider an ${\mathcal{F}}$-regular function $\mathbb{u}:\Omega\times\mathbb{R}\to\mathbb{R}$ such that $\mathbb{u}(\cdot,x)$ is integrable for all $x\in\mathbb{R}$ and let $\Sigma$ be a collection of sub-$\sigma$-algebras of ${\mathcal{F}}$. We define a preference relation on $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ by
$f_{1}\succeq f_{2}\text{ if and only if }\int_{\Omega}\mathbb{u}(\omega,f_{1}(\omega))\mathrm{d}\mathbb{P}(\omega)\geq\int_{\Omega}\mathbb{u}(\omega,f_{2}(\omega))\mathrm{d}\mathbb{P}(\omega).$
For any ${\mathcal{G}}\in\Sigma$ we can directly apply Theorem 8 and find a ${\mathcal{G}}$-regular function $\mathbb{u}_{\mathcal{G}}:\Omega\times\mathbb{R}\to\mathbb{R}$ such that $\mathbb{u}_{\mathcal{G}}(\cdot,x)$ is a version of $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(x)\middle|{\mathcal{G}}\right]$. Now fix $X\in L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$ and let $f$ be any representative element of the equivalence class $X$. Analogously to the argument in Section 5.3, steps 3 and 4, we take a suitable version $h\in\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)\middle|{\mathcal{G}}\right]$ and compute $\Phi_{\mathcal{G}}(\cdot,h(\cdot))$ where $\Phi_{\mathcal{G}}:\Omega\times\mathbb{R}\rightarrow\overline{\mathbb{R}}$ is the right inverse of $\mathbb{u}_{\mathcal{G}}$, that is
Observe that $[\Phi_{\mathcal{G}}(\cdot,h(\cdot))]_{\mathbb{P}}$ does not depend on the choice of the representative $f\in X$, hence the definition $\mathbb{u}_{\mathcal{G}}^{-1}\left(\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(X)\middle|{\mathcal{G}}\right]\right):=[\Phi_{\mathcal{G}}(\cdot,h(\cdot))]_{\mathbb{P}}$ is well posed. In particular, the previous argument formalizes the definition of the mapping ${\mathcal{E}}_{\mathcal{G}}:X\mapsto\mathbb{u}_{\mathcal{G}}^{-1}\left(\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(X)\middle|{\mathcal{G}}\right]\right)$ provided in the statement of Theorem 2. It remains to show that $\left\{{\mathcal{E}}_{\mathcal{G}}\right\}_{{\mathcal{G}}\in\Sigma}$ defined above is a family of time consistent conditional nonlinear expectations such that ${\mathcal{E}}_{0}$ satisfies Assumption 1. It is just a simple checking that for any $X\in L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$, $A\in{\mathcal{G}}$ we have ${\mathcal{E}}_{{\mathcal{G}}}(X\mathbf{1}_{A})={\mathcal{E}}_{{\mathcal{G}}}(X)\mathbf{1}_{A}$. Similarly, ${\mathcal{E}}_{0}({\mathcal{E}}_{{\mathcal{G}}}(X))={\mathcal{E}}_{0}(X)$ for any ${\mathcal{G}}\in\Sigma$ and $X\in L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$. Finally we observe that ${\mathcal{E}}_{0}(X)=\mathbb{u}_{0}^{-1}\left(\int_{\Omega}\mathbb{u}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)\right)$ for $\mathbb{u}_{0}:x\mapsto\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(x)\right]$ and some $f\in X$. Therefore from strict monotonicity and continuity of $\mathbb{u}$ we get: (a) for any $x,y\in\mathbb{R}$, $Z\in L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$ and $A\in{\mathcal{F}}$ with $\mathbb{P}(A)>0$, if $x<y$ then ${\mathcal{E}}_{0}(x\mathbf{1}_{A}+Z\mathbf{1}_{A^{c}})<{\mathcal{E}}_{0}(y\mathbf{1}_{A}+Z\mathbf{1}_{A^{c}})$; (b) for any bounded sequence $X_{n}$ converging $\mathbb{P}$-almost surely to $X$ we have $\lim_{n\to\infty}{\mathcal{E}}_{0}(X_{n})={\mathcal{E}}_{0}(X)$. This concludes the first implication.
Now suppose $\left\{{\mathcal{E}}_{\mathcal{G}}\right\}_{{\mathcal{G}}\in\Sigma}$ is a family of time consistent nonlinear conditional expectations such that ${\mathcal{E}}_{0}$ satisfies Assumption 1. We show that there exists an ${\mathcal{F}}$-regular function $\mathbb{u}:\Omega\times\mathbb{R}\to\mathbb{R}$ such that $\mathbb{u}(\cdot,x)$ is integrable for all $x\in\mathbb{R}$ and, for any $X\in L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$ and ${\mathcal{G}}\in\Sigma$, ${\mathcal{E}}_{\mathcal{G}}$ is of the form
${\mathcal{E}}_{{\mathcal{G}}}(X)=\mathbb{u}_{\mathcal{G}}^{-1}(\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(X)\middle|{\mathcal{G}}\right]),$
with $\mathbb{u}_{\mathcal{G}}(\cdot,x)=\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(x)\middle|{\mathcal{G}}\right]$. For any $f_{1},f_{2}\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ we define the preference order $\succeq$ as follows
$f_{1}\succeq f_{2}\text{ if and only if }{\mathcal{E}}_{0}(X_{1})\geq{\mathcal{E}}_{0}(X_{2}),$
where $X_{1}=[f_{1}]_{\mathbb{P}},X_{2}=[f_{2}]_{\mathbb{P}}\in L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$ are the equivalence classes generated by $f_{1},f_{2}$, and notice that the preference relation $\succeq$ is well defined.
We start by showing that the null sets of the preference as defined in (6) correspond to the null sets of the reference probability measure $\mathbb{P}$ i.e. ${\mathcal{N}}_{\succeq}=\{A\in{\mathcal{F}}\mid\mathbb{P}(A)=0\}$. Let $A\in{\mathcal{F}}$ be such that $\mathbb{P}(A)=0$ and consider $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ with $f$ being a representative element in $X\in L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$. Take any $g\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ being a representative of any $Y\in L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$. Then since $\mathbb{P}(A)=0$, we have that $f\mathbf{1}_{A^{c}}+g\mathbf{1}_{A}$ is still a representative of $X$, thus ${\mathcal{E}}_{0}(X\mathbf{1}_{A^{c}}+Y\mathbf{1}_{A})={\mathcal{E}}_{0}(X)$. This yields $A\in{\mathcal{N}}_{\succeq}$. Now suppose $A\in{\mathcal{F}}$ is such that $\mathbb{P}(A)>0$ and consider $x,y\in\mathbb{R}$ with $x<y$. For any choice of $Z\in L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$, monotonicity of ${\mathcal{E}}_{0}$ implies that
${\mathcal{E}}_{0}(x\mathbf{1}_{A}+Z\mathbf{1}_{A^{c}})<{\mathcal{E}}_{0}(y\mathbf{1}_{A}+Z\mathbf{1}_{A^{c}}),$
which in turn yields $A\notin{\mathcal{N}}_{\succeq}$.
Secondly, we observe that $\succeq$ satisfies (SM) and (PC) as an immediate consequence of Assumption 1. Finally, notice that for any $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ the functional $T(f)={\mathcal{E}}_{0}(X)$, with $X=[f]_{\mathbb{P}}\in L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$, represents $\succeq$. By the properties of the family $\left\{{\mathcal{E}}_{\mathcal{G}}\right\}_{{\mathcal{G}}\in\Sigma}$, for any sub-$\sigma$-algebra ${\mathcal{G}}$ we have ${\mathcal{E}}_{0}(X\mathbf{1}_{A})={\mathcal{E}}_{0}({\mathcal{E}}_{{\mathcal{G}}}(X)\mathbf{1}_{A})$ so that $T(f\mathbf{1}_{A})=T(g\mathbf{1}_{A})$ for any $A\in{\mathcal{G}}$ and $g\in{\mathcal{E}}_{{\mathcal{G}}}(X)$. Hence $T$ is conditionable and in particular, by Theorem 4, $\succeq$ satisfies (ST). By Theorem 7 we can find a measure $\mathbb{Q}$ and a state-dependent utility $\tilde{\mathbb{u}}$ such that Eq. (11) holds, adopting the pair $(\tilde{\mathbb{u}},\mathbb{Q})$. We observe that $\mathbb{Q}\sim\mathbb{P}$ and hence we can redefine $\mathbb{u}(\omega,x)=\frac{d\mathbb{Q}(\omega)}{d\mathbb{P}}\tilde{\mathbb{u}}(\omega,x)$, where $\frac{d\mathbb{Q}}{d\mathbb{P}}$ is a version of the Radon Nikodym derivative, and observe that Eq. (11) still holds true for ($\mathbb{u},\mathbb{P}$), with $\mathbb{u}$ being ${\mathcal{F}}$-regular. Applying Theorem 8 we indeed can define a ${\mathcal{G}}$-regular function $\mathbb{u}_{\mathcal{G}}:\Omega\times\mathbb{R}\to\mathbb{R}$, with $\mathbb{u}_{\mathcal{G}}(\cdot,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{G}},\mathbb{P})$ for every $x\in\mathbb{R}$, as a version of $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(x)\middle|{\mathcal{G}}\right]$ for every $x\in\mathbb{R}$. In this way, for an appropriate choice of $h\in\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)\middle|{\mathcal{G}}\right]$, $\Phi_{\mathcal{G}}(\cdot,h(\cdot))\in\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ where $\Phi_{\mathcal{G}}:\Omega\times\mathbb{R}\rightarrow\left[-\infty,+\infty\right]$ defined by
$\Phi_{\mathcal{G}}(\omega,x):=\inf\left\{y\in\mathbb{R}\middle|\mathbb{u}_{\mathcal{G}}(\omega,y)>x\right\}$
is ${\mathcal{G}}\otimes\mathbb{B}_{\mathbb{R}}$ measurable. Note that the map $X\mapsto[\Phi_{\mathcal{G}}(\cdot,h(\cdot))]_{\mathbb{P}}$ is well defined, since the procedure above does not depend (up to $\mathbb{P}$-null events) on the particular choice of a representative of $X$, and we shall write $\mathbb{u}_{\mathcal{G}}^{-1}\left(\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(X)\middle|{\mathcal{G}}\right]\right)=[\Phi_{\mathcal{G}}(\cdot,h(\cdot))]_{\mathbb{P}}$. By the previous argument any element $g\in{\mathcal{E}}_{{\mathcal{G}}}(X)$ is also an element of $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ and therefore $\mathbb{u}_{\mathcal{G}}^{-1}\left(\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(X)\middle|{\mathcal{G}}\right]\right)={\mathcal{E}}_{{\mathcal{G}}}(X)$.
Remark 17.
It is worthy of note that the “taking out what is known” property in Definition 1 jointly with Assumption 1 imply that for any $Y\in L^{\infty}(\Omega,\mathcal{G},\mathbb{P})$ we have ${\mathcal{E}}_{\mathcal{G}}(Y)=Y$. To see this, first observe that by strict monotonicity and time consistency one can show ${\mathcal{E}}_{\mathcal{G}}(x)=x$ for $x\in\mathbb{R}$. Consequently, applying the “taking out what is known” property, it follows that ${\mathcal{E}}_{\mathcal{G}}(Z)=Z$ for any ${\mathcal{G}}$-measurable simple function $Z$. To conclude, for a general $Y\in L^{\infty}(\Omega,\mathcal{G},\mathbb{P})$ we can choose a sequence of ${\mathcal{G}}$-measurable simple functions $(\hat{Y}_{N})_{N}$ such that $\hat{Y}_{N}\to_{N}Y$ with respect to the sup norm, then it can be shown that $||{\mathcal{E}}_{\mathcal{G}}(Y)-\hat{Y}_{N}||_{\infty}\to_{N}0$ which in turn yields ${\mathcal{E}}_{\mathcal{G}}(Y)=Y$.
@@ Appendix A Appendix: auxiliary results
Lemma 18.
Suppose $\succeq$ satisfies (SM) and (PC), then $\succeq$ is monotone on $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, that is, for $f,g\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, $f(\omega)\geq g(\omega)$ for all $\omega\in\Omega$ implies $f\succeq g$.
Proof.
We begin by showing the result for simple functions. Consider two simple functions $f,g\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ such that $f(\omega)\geq g(\omega)$ for all $\omega\in\Omega$. It is possible to find a common partition $\pi:=\left\{A_{1},A_{2},...,A_{N}\right\}\subseteq{\mathcal{F}}$ for some $N\in\mathbb{N}$ such that we can write $f=\sum_{i=1}^{N}x_{i}\mathbf{1}_{A_{i}}$, $g=\sum_{i=1}^{N}y_{i}\mathbf{1}_{A_{i}}$, with $x_{i},y_{i}\in\mathbb{R}$ for $i=1,...,N$. We have
$f=\sum_{i=1}^{N}x_{i}\mathbf{1}_{A_{i}}\succeq y_{1}\mathbf{1}_{A_{1}}+\sum_{i=2}^{N}x_{i}\mathbf{1}_{A_{i}}\succeq y_{1}\mathbf{1}_{A_{1}}+y_{2}\mathbf{1}_{A_{2}}+\sum_{i=3}^{N}x_{i}\mathbf{1}_{A_{i}}\succeq\ldots\succeq\sum_{i=1}^{N}y_{i}\mathbf{1}_{A_{i}}=g,$
where $\succeq$ will be $\sim$ if $x_{i}=y_{i}$ and $\succ$ if $x_{i}>y_{i}$ by (SM). We conclude $f\succeq g$ for any pair of simple functions.
Now consider $f,g\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ such that $f(\omega)\geq g(\omega)$ for all $\omega\in\Omega$. Let $\left(f_{n}\right)_{n},\left(g_{n}\right)_{n}\subseteq\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ be sequences of simple functions such that $f_{n}\downarrow_{n}f$ and $g_{n}\uparrow_{n}g$ so that $f_{n}\geq g_{n}$ for every $n\in\mathbb{N}$. Therefore $f_{n}\succeq g_{n}$ from the previous point, for every $n\in\mathbb{N}$. Now assume by contradiction that $g\succ f$, then applying (PC) we find $n_{1}$ such that $g_{n_{1}}\succ f$ and applying again (PC) we find $n_{2}$ such that $g_{n_{1}}\succ f_{n_{2}}$. By monotonicity on simple functions we have $g_{n}\succeq g_{n_{1}}\succ f_{n_{2}}\succeq f_{n}$ for every $n\geq n_{1}\vee n_{2}$ and hence a contradiction. ∎
Lemma 19.
@@
Proof.
We prove Item (i) in the spirit of [15] (Proposition 1.1.13 and Remark 1.1.14) as follows. Let $m>0$. For $n\geq 1$, $k=0,1,...,2\cdot 2^{n}-1$, we define
@@
The newly constructed process is piecewise constant over intervals of length $\frac{m}{2^{n}}$ and right continuous. The map $(\omega,s)\mapsto u^{(n)}(\omega,s)$ is ${\mathcal{F}}\otimes\mathbb{B}_{\mathbb{R}}$-measurable. By right continuity of $u$ we have that $u^{(n)}(\omega,s)\to u(\omega,s)$ for every pair $(\omega,s)\in\Omega\times[-m,m]$ as $n\to\infty$. Therefore we obtain that the map $(\omega,s)\mapsto u(\omega,s)$ is itself ${\mathcal{F}}\otimes\mathbb{B}_{[-m,m]}$-measurable as pointwise limit of a sequence of measurable functions. Finally, by letting $m\to\infty$ we have that $(\omega,s)\mapsto u(\omega,s)$ is ${\mathcal{F}}\otimes\mathbb{B}_{\mathbb{R}}$-measurable. By the latter argument we have that $\omega\mapsto u(\omega,x)$ is ${\mathcal{F}}\otimes\mathbb{B}_{\mathbb{R}}-\mathbb{B}_{\mathbb{R}}$ measurable. Moreover, for any ${\mathcal{F}}-\mathbb{B}_{\mathbb{R}}$ measurable and bounded $f$, the map $\varphi:\omega\mapsto(\omega,f(\omega))$ is ${\mathcal{F}}-{\mathcal{F}}\otimes\mathbb{B}_{\mathbb{R}}$ measurable (see [1], Lemma 4.49). Hence, the composition $(u\circ\varphi):\omega\mapsto u(\omega,f(\omega))$ is ${\mathcal{F}}-\mathbb{B}_{\mathbb{R}}$ measurable. To conclude, since $u$ is nondecreasing, ${\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})\ni u(\omega,-\lVert f\rVert_{\infty})\leq u(\omega,f(\omega))\leq u(\omega,\lVert f\rVert_{\infty})\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ thus concluding the proof of Item (ii). ∎
Proposition 20.
@@
$A_{\text{cont}}:=\{\omega\in\Omega\mid x\mapsto u(\omega,x)\text{ is continuous}\}\in{\mathcal{F}}$
and $\mathbb{P}(A_{\text{cont}})=1$.
Proof.
In the present proof we shall rely on few concepts related to the theory of stochastic processes and in particular on stopping times. We refer the reader to [15], page 6, for an exhaustive treatment of the notions involved.
Lemma 19 guarantees that $T_{u}:\mathcal{L}^{\infty}(\Omega,{\mathcal{F}})\to\mathbb{R}$ is well defined and finite valued. Fix $M>0$ and define $X_{t}(\omega):=u(\omega,t-M)$ for $t\geq 0$, omitting the dependence on $M$ for simplicity. Take also the filtration ${\mathcal{F}}_{t}:={\mathcal{F}},t\geq 0$, which is trivially right continuous by definition: $\bigcap_{s>t}{\mathcal{F}}_{s}={\mathcal{F}}={\mathcal{F}}_{t}$. Then $(X_{t})_{t}$ is an RCLL process, adapted to $({\mathcal{F}}_{t})_{t}$. Set $\Delta X_{0}=0$ and for $t>0$ set $\Delta X_{t}:=X_{t}-\sup_{0\leq s<t}X_{s}=X_{t}-\lim_{s\uparrow t}X_{s}$.
Following the idea in [22], we define $\tau^{\varepsilon}$ as the first jump time for the size $\varepsilon>0$, namely
@@
where equality ()($\star$)( ⋆ ) follows from posing $\Delta X_{0}=0$. Looking back at $u(\omega,\cdot)$, $\tau^{\varepsilon}(\omega)$ is the first time $u(\omega,\cdot)$ jumps, in the interval $(-M,+\infty)$, with a jump size strictly greater than $\varepsilon$, with the usual convention that such time is set to be +$\infty$+ ∞ if no such jump actually occurs. By [22], $\tau^{\varepsilon}:\Omega\rightarrow[0,+\infty]$ is a $({\mathcal{F}}_{t})_{t}$-stopping time, hence in particular it is ${\mathcal{F}}$-measurable by definition of such a filtration. Setting
@@
Take now $s_{k}:=(Z-\frac{1}{k}\mathbf{1}_{B})^{+}$ for $k\in\mathbb{N}$. $s_{k}$ is ${\mathcal{F}}-\mathbb{B}_{[0,+\infty)}$ measurable and bounded. Observe that since $(X_{t})_{t}$ is RCLL, by [15] Remark 1.1.14 we have $(t,\omega)\mapsto X_{t}(\omega)$ is $\mathbb{B}_{[0,+\infty)}\otimes{\mathcal{F}}-\mathbb{B}_{[0,+\infty)}$ measurable, and $\omega\mapsto(\omega,Y(\omega))$ is ${\mathcal{F}}-{\mathcal{F}}\otimes\mathbb{B}_{[0,+\infty)}$ measurable for every ${\mathcal{F}}-\mathbb{B}_{[0,+\infty)}$ measurable bounded $Y:\Omega\rightarrow[0,+\infty)$. Hence $\omega\mapsto X_{s_{k}(\omega)}(\omega)$ is ${\mathcal{F}}-\mathbb{B}_{[0,+\infty)}$ measurable. By continuity from below, argued at the beginning of this proof,
@@
Corollary 21.
Under the same assumptions of Proposition 20, there exists a function $\widehat{u}$ such that $\widehat{u}(\omega,\cdot)$ is nondecreasing and continuous for every $\omega\in\Omega$, $E=\{\omega\in\Omega\mid u(\omega,x)=\widehat{u}(\omega,x)\,\forall x\in\mathbb{R}\}\in{\mathcal{F}}$ and ${\mathbb{P}}(E)=1$. In particular, $T_{u}(f)=T_{\widehat{u}}(f),\,\forall f\in\mathcal{L}^{\infty}(\Omega,{\mathcal{F}})$. Finally, if $u(\omega,\cdot)$ is strictly increasing on $\mathbb{R}$ for every $\omega\in\Omega$, $\widehat{u}(\omega,\cdot)$ can be taken strictly increasing on $\mathbb{R}$ for every $\omega\in\Omega$.
Proof.
As a consequence of Proposition 20, $A_{cont}\in{\mathcal{F}}$. Set $A:=A_{cont}$ and define now $\widehat{u}(\omega,x):=u(\omega,x)\mathbf{1}_{A}(\omega)+x\mathbf{1}_{\Omega\setminus A}$. Then clearly $\widehat{u}(\omega,\cdot)$ is continuous and nondecreasing either $\omega\in A$ or $\omega\in\Omega\setminus A$. Further for any fixed $x\in\mathbb{R}$ the map $\omega\mapsto\widehat{u}(\omega,x)$ is ${\mathcal{F}}$ measurable. As $u(\omega,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ implies $\widehat{u}(\omega,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ for every $x\in\mathbb{R}$ we can invoke Lemma 19 to obtain that $\widehat{u}$ is ${\mathcal{F}}\otimes\mathbb{B}_{\mathbb{R}}-\mathbb{B}_{\mathbb{R}}$ measurable, $T_{\widehat{u}}$ is well defined in ${\mathcal{L}}^{\infty}(\Omega,{\mathcal{F}})$ and $T_{u}\equiv T_{\widehat{u}}$.